• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Glashausser

    Log entry time 16:05:22 on August10,2003

    Entry number 110055

    keyword=Shift Summary 8/10/03

    8:00 Starting Kin 8.15
    10:10 Beam HWP inserted
    10:30 See Halog entry (and run sheet for KIN 8.15) for comments on calculation of total CODA counts needed. No background correction necessary. VS halog
    entry (110016) discussing pion cut doesn't affect this calculation either.
    10:57 Beam HWP removed.
    Change target fields to 270 deg.
    11:42 Finish KIN 8.15
    11:43 Changing to KIN 8.16 (p=1.853)
    Start off with factor of 128.2 for septum.
    Setting up NMR and EPR polar measurements
    12:00 Starting NMR measurement
    12:12 NMR result: 36.83%
    EPR result: 35.88%
    12:27 Beam back--start 3407 (C foil)
    13:00 CHL crash
    15:00 MCC says it will be four hours from the time they fix the problem. The problem seems to be a compressor bearing. Kent P suggested that we ramp down the septum, which we are now doing. Scott Spiegel suggested we keep other magnets as is unless they start to trip.